OpenStudy (anonymous):

What is the equation of the tangent line passing through the point (0, 3) of the graph of the function f(x) = x^2 − 2x + 3? y = –2x + 3 y = 3x + 1 y = 2x − 3 y = –3x + 3

OpenStudy (michele_laino):

here we have to apply this formula: \[\Large y - 3 = f'\left( 0 \right)\left( {x - 0} \right)\]

OpenStudy (michele_laino):

where f ' (0) is the first derivative of your function at point x=0